OpenStudy (anonymous):

MEDAL! F = 9 over 5C + 32 f = 10

OpenStudy (anonymous):

?

OpenStudy (ayyookyndall):

the order of operations says you must subtract the 32 from both sides first, if you want to solve for C: F = (9/5)C + 32 F - 32 = (9/5)C (5/9)(F - 32) = C

OpenStudy (anonymous):

oh, okay.

OpenStudy (anonymous):